OpenStudy (anonymous):

A juggler is performing an act by juggling several balls. The juggler throws the balls up at an initial height of 3.9 feet with a speed of 14.8 feet per second. If the juggler did not catch a ball, about how long will it take the ball to hit the floor? a) 1.45 secs b) 1.14 secs c) 0.28 secs d) 0.21 secs

OpenStudy (anonymous):

I only know SI units, so this may be wrong but the problem is v = u +at where a is negative gravity so 0 = 14.8 - 32.174 t so t = 0.46 seconds. so how high did it get? s = ut +1/2 at^2 s = 14.8(0.46) +0.5 * -32.174*0.46^2 s = 3.4 so It started at 3.9 feet and went another 3.4 so total height is 7.3 s = ut +1/2 at^2 7.3 = 0 + 0.5 *32.174 *t^2 t = 0.67 so total t = 0.46 +0.67 = 1.13 seconds

OpenStudy (anonymous):

can u answer this diarmuid wrenne Explain at a molecular level what happens during latent heat of fusion or evaporation

OpenStudy (anonymous):

sorry. I'm not much good at the theory stuff!!